OpenStudy (trojanpoem):

If f'(x) = Ax^2 + b (b,a are constants) f(0) + f'(0) = 1 , f(1) = 3 then , f'(1) = .....

OpenStudy (trojanpoem):

I fail to find b :/ the final answer is 4

OpenStudy (anonymous):

b = 1

OpenStudy (trojanpoem):

Show me your work.

OpenStudy (anonymous):

no!)

OpenStudy (anonymous):

i haven't do anything, its only my thoughts

OpenStudy (anonymous):

f'(x) = 2Ax

OpenStudy (trojanpoem):

f'(x) = Ax^2 + b.

OpenStudy (anonymous):

and f'(0)= 2a*0

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

just a moment

OpenStudy (trojanpoem):

It's already the first derivative.Take your time.

OpenStudy (perl):

did you take the integral?

OpenStudy (perl):

f'(x) =ax^2 + b f(0) + f'(0) = 1 , f(1) = 3 f(x) = ax^3/3 + bx + c f(0) = a(0) + b(0) + c f ' (0) = a(0) + b

OpenStudy (trojanpoem):

I did , The result was 2 unknowns with 1 equation. A = 6

OpenStudy (trojanpoem):

c + b = 1 (1) f(1) = 1/3 * a + b + c = 3 from (1) 1 + 1/3 a = 3 so a = 6

OpenStudy (perl):

correct

OpenStudy (trojanpoem):

f'(1) = a + b = 6 + b. The problem now is to find another equation to find b.

OpenStudy (trojanpoem):

Clarify a bit, I can't understand, well.

OpenStudy (perl):

so we found that a = 6. lets plug back into original f(x) and f ' (x) f ' (x) = 6x^2 + b f (x) = 6/3 x^3 + bx + c = 2x^3 + bx + c

OpenStudy (trojanpoem):

Okie , now ?

OpenStudy (perl):

hmm, there might be a typo in the directions, im not getting b = 4 can you upload the original directions, take a screen shot

OpenStudy (trojanpoem):

b =-2 a + b = 4

OpenStudy (trojanpoem):

The book is in front of me. I didn't mistype it. "No PC cable or PC Bluetooth"

OpenStudy (perl):

These are the directions: If f'(x) = Ax^2 + b (b,a are constants) f(0) + f'(0) = 1 f(1) = 3 find f ' (1)

OpenStudy (trojanpoem):

That's it.

OpenStudy (trojanpoem):

either the book writer was mistaken or there is magic in it.

OpenStudy (perl):

right, either its a typo or something else. what is the name of this book, maybe i can check google book results, the question might pop up

OpenStudy (trojanpoem):

Sorry , I was afk.